# Insulator-Metal Transition and Magnetic Crossover in Bilayer Graphene

### v2

@@ Abstract
Introduction. The effects of perpendicular magnetic fields on the electronic properties of monolayer graphene have been studied in a variety of contexts over the years. Landau levels[1], integer [2] and fractional[3] quantum Hall effects were demonstrated experimentally within several years of graphene discovery[4]. Very recently, a giant magnetoresistance of Dirac electrons in high-mobility graphene has been discovered [5]. In addition, it has been shown that the orbital magnetic response of graphene can be controlled by modifying the number of charge carriers via gating. Perfect diamagnetism at the charge neutrality point [6] is superseded by paramagnetism as the Fermi level approaches a van Hove singularity in the density of states [7] making it possible to switch the monolayer between diamagnetic and paramagnetic regimes[8]. All of these effects have been experimentally observed [9] and thoroughly investigated theoretically [10, 11, 12].
In recent years, bilayer and multilayer graphene have garnered significant attention following the discovery of flat bands in twisted bilayer graphene that host superconductivity [13]. These structures are also remarkable for their orbital magnetic properties [14, 10]. A key feature of multilayer structures is their sensitivity to in-plane magnetic fields, which are coupled to the in-plane component of the orbital magnetic moment due to the looping motion of the electron between layers [15, 16, 17, 18, 19]. In a bilayer system, for instance, the vector potential associated with the in-plane magnetic field can have opposite signs in the two layers, while remaining constant in each layer. This causes the band structures of the layers (before turning on interlayer coupling) to shift in opposite directions in momentum space, leading to non-trivial modifications of the band structure when interlayer coupling is considered. This effect was experimentally demonstrated a few years ago in twisted bilayer graphene, where an in-plane magnetic field was used to alter the separation of nearby Dirac cones[20].
@@
Numerically calculated $\chi$ as a function of $\Phi$ at $\mu=0$ is shown in Fig. 2 for several values of $V$. The most important feature exhibited by the plots is the sharp transition from diamagnetic to paramagnetic behavior at the point of IM transition. It is possible to understand the origins of this behavior by considering individual terms in Eq. (6) for $T\rightarrow 0$. The first term is negative, i.e., diamagnetic, and has a very weak dependence over the chemical potential. The second and third terms are both positive and therefore paramagnetic. Specifically, the second term describes inter-band paramagnetism [33] and the final term is a Fermi contour contribution that arises from the direct coupling of the in-plane magnetic field with the fluctuating orbital moment ([27]): this term vanishes when the system is an insulator.
@@
Giovanni Vignale January 15, 2025

### v3

@@ Abstract
Introduction. The effects of perpendicular magnetic fields on the electronic properties of monolayer graphene have been studied in a variety of contexts over the years. Landau levels[1], integer [2] and fractional[3] quantum Hall effects were demonstrated experimentally within several years of graphene discovery[4]. Very recently, a giant magnetoresistance of Dirac electrons in high-mobility graphene has been discovered [5]. In addition, it has been shown that the orbital magnetic response of graphene can be controlled by modifying the number of charge carriers via gating. Perfect diamagnetism at the charge neutrality point [6, 7, 8] is superseded by paramagnetism as the Fermi level approaches a van Hove singularity in the density of states [9] making it possible to switch the monolayer between diamagnetic and paramagnetic regimes[10]. All of these effects have been experimentally observed [11] and thoroughly investigated theoretically [12, 13, 14].
In recent years, bilayer and multilayer graphene have attracted significant interest due to their highly tunable properties under mechanical strain, twisting, and external fields. Their remarkable orbital magnetic properties [15, 12] and the discovery of flat bands in twisted bilayer graphene, which host superconductivity [16], further highlight their importance. A recent study on Bernal-stacked bilayer graphene reveals a complex multi-cone band structure under finely tuned displacement fields [17]. Another important property of this system is its strong response to in-plane magnetic fields, which couple to the in-plane component of the orbital magnetic moment through the interlayer looping motion of electrons [18, 19, 20, 21, 22]. In a bilayer system, for instance, the vector potential associated with the in-plane magnetic field can have opposite signs in the two layers, while remaining constant in each layer. This causes the band structures of the layers (before turning on interlayer coupling) to shift in opposite directions in momentum space, leading to non-trivial modifications of the band structure when interlayer coupling is considered. This effect was experimentally demonstrated a few years ago in twisted bilayer graphene, where an in-plane magnetic field was used to alter the separation of nearby Dirac cones[23].
@@
Notice that the 2D susceptibility has the dimensions of a length. To give a clearer idea of its magnitude we divide the numerically calculated $\chi_{2D}$ by the absolute value of the out-of-plane orbital susceptibility of a charge-neutral graphene monolayer at $T=300$K [7, 6, 8] i.e., $|\chi_{orb}|=\frac{1}{6\pi}\frac{e^{2}v^{2}}{c^{2}}\frac{1}{k_{B}T}\approx 2.7\times 10^{-14}m$. The ratio $\frac{\chi_{2D}}{|\chi_{orb}|}$ is plotted as a function of $\Phi$ at $\mu=0$ in Fig. 2 for several values of $V$. The most striking feature exhibited by the plots is the sharp transition from diamagnetic to paramagnetic behavior at the point of IM transition. It is possible to understand the origins of this behavior by considering individual terms in Eq. (6) for $T\rightarrow 0$. The first term is negative, i.e., diamagnetic, and has a very weak dependence over the chemical potential. The second and third terms are both positive and therefore paramagnetic. Specifically, the second term describes inter-band paramagnetism [36] and the final term is a Fermi contour contribution that arises from the direct coupling of the in-plane magnetic field with the fluctuating orbital moment ([30]): this term vanishes when the system is an insulator.
@@
Giovanni Vignale February 19, 2025